OpenStudy (anonymous):

( cos theta - cot theta ) / 1- sin theta = - cot theta prove the identity.. i'm having troubling making the LS equal the RS

OpenStudy (anonymous):

hmm, i'll try not to draw this time :D

hero (hero):

lol

OpenStudy (anonymous):

(cosθ-cotθ)/1-sin θ = - cot θ Work with left side (cosθ- cosθ/sinθ) / 1-sinθ (cos/1 - cosθ/sinθ) / 1-sinθ (cosθsinθ-cosθ)/sinθ / 1-sinθ cosθ (sinθ-1)/ sinθ / 1-sinθ Now at this part, cosθ become cotθ because cosθ/sinθ = cotθ cotθ( sinθ-1) / 1-sinθ (sinθ-1) and 1-sinθ cancel out left with a -cotθ = -cotθ Hope it's not too confusing :|

hero (hero):

Looks worse than the drawings

OpenStudy (anonymous):

i take that as a compliment :X

OpenStudy (anonymous):

cos(a)-cot(a)/1-sin(a) =[sin(a)cos(a)-cos(a)]/sin(a)(1-sin(a)) =cos(a)(sin(a)-1)/-[sin(a)(sin(a)-1] =-cot(a)

hero (hero):

\(\LaTeX\) is very easy to learn

OpenStudy (anonymous):

where

hero (hero):

type `\(\LaTeX\)` to produce this: \(\LaTeX\)

hero (hero):

type `\(\frac{2x}{5}\)` to produce this: \(\frac{2x}{5}\)

OpenStudy (anonymous):

\(\LaTeX\)

hero (hero):

Congratulations. That's the equivalent of learning how to print "Hello World" in java

OpenStudy (anonymous):

cool! thx :D

OpenStudy (anonymous):

@soapia it helps if you write out what i did on paper, so u can see it or maybe @Hero can \(\LaTeX\) for u

hero (hero):

@Shido88, you can \(\LaTeX\) All you do is just put everything you type between `\( \)` Think of them as quotation marks for math

OpenStudy (anonymous):

(cosθsinθ-cosθ)/sinθ / 1-sinθ cosθ (sinθ-1)/ sinθ / 1-sinθ ... you cross multiplied the numerator..?

hero (hero):

There are some special symbols though. For example, to do fractions, you have to use \frac{a}{b}, where a = numerator expression, b = denominator expression

OpenStudy (anonymous):

whoops.. i mean this step (cos/1 - cosθ/sinθ) / 1-sinθ to this step.. (cosθsinθ-cosθ)/sinθ / 1-sinθ u cross multiplied the numerator?

hero (hero):

So typing `\(\frac{a}{b}\)` produces: \(\frac{a}{b}\)

OpenStudy (anonymous):

@soapia no, I factor out the cosθ

OpenStudy (anonymous):

\(Shido\) oh ok

hero (hero):

`\(2x^2\)` produces \(2x^2\)

hero (hero):

`\(2 \times 5\)` produces \(2 \times 5\)

hero (hero):

`\(16 \div 4 = 4\)` produces \(16 \div 4 = 4\)

hero (hero):

`\(\cos\theta\)` produces \(\cos\theta\)

hero (hero):

`\(x = \frac{b \pm \sqrt{b^2 - 4ac}}{2a}\)` produces \(x = \frac{b \pm \sqrt{b^2 - 4ac}}{2a}\)

hero (hero):

You only use the \ symbol before tag words like \frac \cos \times \div \sqrt, etc...

hero (hero):

And each line of algebraic expressions go between `\( \)`

OpenStudy (anonymous):

ahh computer froze for a while after doing Latex! @soapia sorry for late reply, you don't exactly multiply them cuz this is a minus in between. When it's a multiplication or division in between then you can but this is minus, so just basictly combine the top part and bottom part. Bottom part is the same so understood to be itself

OpenStudy (anonymous):

@Hero where did u learn all this?

hero (hero):

A lot of users know how to use \(\LaTeX\). there's a whole group dedicated to it. There are books written on it. It's been out since the 60's or 70's I believe.

hero (hero):

No, that doesn't imply that I've been around for that long.

hero (hero):

http://en.wikipedia.org/wiki/LaTeX

OpenStudy (anonymous):

lol, i know lots of people knows how to use it but it just seems amazing that you can remember to do all that. Maybe I'm just new to it i guess :)

OpenStudy (anonymous):

haaa! found a youtube video lol !

hero (hero):

Be careful, there are different quotation marks that are used. Here, we use `\( \)` or `\[ \]`

hero (hero):

Others use `& &` or `$ $` or `; ;`

OpenStudy (anonymous):

Cool! will try next time

hero (hero):

The first one means "begin Latex" the second one means "end Latex"

OpenStudy (anonymous):

I see